OpenStudy (anonymous):

Jodie bicycles 7 km/h faster than Praedup. In the same time it takes Praedup to bicycle 39 km, Jodie can bicycle 60 km. How fast does each bicyclist travel?

OpenStudy (anonymous):

Speed of Pradup=x Jodie bike=x+7 (x)t=39 (x+7)t=60

OpenStudy (anonymous):

8

OpenStudy (anonymous):

t=60/(x+7) Plug that back into first equation x* 60/(x+7)=39